### v2

# Attosecond Core-level Photoionization Delays of Aromatic Molecules

Attosecond Core-level Photoionization Delays
of Aromatic Molecules
Attosecond photoionization delays are a unique probe of the structure and the electronic dynamics of matter, but valence-shell photoionization dynamics of polyatomic systems are often complicated and challenging to interpret because of the delocalized nature of the electronic wave functions. Here we use attosecond soft-X-ray pulses from LCLS to probe element-specific core-level photoionization delays between N-1s and C-1s core shells of aromatic azabenzene molecules containing one to three nitrogen atoms. Using a self-referenced technique, we measure absolute N-1s photoionization delays of up to 300 as, which increase with the number of nitrogen atoms in the molecule. Comparisons to molecular photoionization calculations reveal the dominant mechanisms to be the increasing trapping of the photoelectron wavefunction in the molecular plane, as well as the symmetry-induced decrease of the coupling strength among the partial waves. This study demonstrates the unique opportunities opened by measurements of core-level photoionization delays for unravelling attosecond electronic dynamics in the building blocks of biomolecules and molecular optoelectronics.
Photoionization or photoemission is often conceptualized as a half-scattering event between a photoelectron wave and matter (?, ?). Owing to the atomic length scale of the electron de-Broglie wave, it is a sensitive probe of the atomic structure of matter (?). Recently, attosecond metrology has added the unique aspect of temporal resolution on atomic time scales, which is absent from traditional collision physics. This advance (?, ?), which has been highlighted in the 2023 Nobel Prize in Physics (?), has integrated time-domain access to multi-electron dynamics into the powerful framework of electron-scattering physics. Not unexpectedly, these unique opportunities have created a flourishing research field with applications to molecules (?, ?, ?, ?, ?, ?), clusters (?, ?), liquids (?) and solids (?, ?, ?, ?). One conclusion from these recent studies is that photoionization delays are sensitive to the spatial extension of electronic wave functions (?). Whereas this property offers exciting avenues for studying the dynamics of electron-hole (de)localization dynamics (?), it may obscure other mechanisms governing photoionization delays because of the strong spatial delocalization characteristic of the valence shell. Moreover, the delocalized nature of valence-shell wavefunctions prevents atomic-scale resolution of the photoionization event and the spectral congestion limits the complexity of the systems that can be studied (?).
Here, we demonstrate the unique opportunities that arise by transposing the research field of attosecond chronoscopy (?) from the extreme-ultraviolet to the X-ray domain. This adds element specificity and site selectivity to the method, while simultaneously confining the emission site of the photoelectron wave to a point-like source. Moreover, core-shell photoelectron spectra have a much simpler structure than valence-shell spectra because they feature a single band per atom of each type (when neglecting satellite lines), which facilitates the extension of the method to complex samples. This advance is made possible by groundbreaking progress in free-electron laser (FEL) science that has led to the generation of attosecond pulses in the X-ray domain through the XLEAP technique (?). The long-standing challenge of temporal jitter between the FEL pulses and laser pulses is solved by using attosecond angular streaking with a circularly-polarized infrared laser pulse and detecting the photoelectron momentum distribution in the polarization plane (?, ?).
The present experiments aim at a fundamental understanding of the attosecond dynamics of electron trapping in aromatic systems (i.e., benzene derivatives), specifically the multiple scattering phenomena that lead to transient trapping of photoelectrons following their release from specific atoms. Our results provide answers to important questions on mechanisms that regulate electron trapping, and the duration over which electrons are trapped. For this research, we chose the azabenzene molecules pyridine, pyrimidine and -triazine, depicted in Fig. 1A. We produced attosecond X-ray pulses with energies just above the nitrogen K-edge located at 405 eV and measured the photoionization delays between the electrons emitted from the nitrogen -shell (N-1s) and those emitted from the carbon -shell (C-1s). Since the latter have kinetic energies of more than 110 eV, the C-1s electrons are emitted with negligible photoionization delays ( as), defining a self-referenced technique for accessing absolute photoionization delays of the N-1s electrons. We measure N-1s photoionization delays of up to 300 as close to the nitrogen -shell threshold, which generally decrease with increasing kinetic energy but display local maxima at 3-15 eV above the threshold. The measured delays also markedly increase from pyridine to pyrimidine and -triazine. These experimental results are compared with state-of-the-art core-level photoionization calculations, which reveal the presence of several shape resonances in this low kinetic-energy region. The comparison of experiment and theory reveals that the photoelectrons originating from the nitrogen -shell are trapped with increasing efficiency in the molecular plane in the sequence pyridine, pyrimidine, -triazine, i.e., with increasing number of nitrogen atoms. Additionally, the increasing symmetry of the molecular scaffold (from C in pyridine to D in -triazine) reduces the coupling among the photoelectron partial waves, which frustrates the decay of the trapped partial waves with high angular momentum. Our study reveals two important mechanisms regulating electronic dynamics in aromatic systems, i.e., the dependence of photoelectron trapping on molecular structure and molecular symmetry. Pyrimidine is a building block of the nucleobases thymine, cytosine and uracil. All three azabenzene molecules are widespread structural motifs in biomolecules, drugs and molecular optoelectronics (?, ?, ?, ?). The discovered mechanisms therefore suggest guidelines for understanding electronic dynamics in biomolecules and molecular optoelectronics.
Figure 1 illustrates the principle of the present experiments, carried out at the time-resolved atomic, molecular, and optical science (TMO) experimental hutch (?) of the Linac Coherent Light Source (LCLS). Linearly polarized attosecond X-ray pulses with a tunable central energy were obtained using the XLEAP technique (?) and were superimposed with circularly polarized short-wave infrared (IR, 1.3 m, 40 fs) pulses in the interaction region of a co-axial velocity-map-imaging (c-VMI) spectrometer. A continuous molecular beam containing pyridine, pyrimidine or -triazine intersected with the overlapping X-ray and IR pulses in the interaction region. The photoelectrons created from nitrogen -shell ionization of the sample were streaked by the IR pulses and projected onto a position-sensitive imaging detector with a central hole. A photoelectron image recorded by ionizing pyrimidine with X-ray pulses centered at 454 eV is shown in Fig. 1A, whereby the inner (outer) ring corresponds to ionization of the nitrogen (carbon) -shell electrons. Figures 1B, 1C, and 1D show how the relative photoionization delays between the N-1s and C-1s photoelectrons were determined. Because each photoelectron was angularly streaked into the direction of the IR-laser vector potential at the instant of photoionization (inset of Fig.1A), the relative streaking angle between N-1s and C-1s electrons directly encodes their relative photoionization delay. The conversion from relative streaking angle to the relative photoionization delay is simply given by , where is the period of the streaking laser (?, ?, ?, ?, ?, ?, ?).
Figure 2 illustrates how the relative streaking angles were obtained from the experimental data. Due to the large temporal jitter (500 fs) between the arrival time of XFEL and IR laser pulses (?), the streaking directions of the core-shell photoelectrons were completely random from shot to shot. The information of relative photoionization time delay is therefore only encoded in the relative streaking angle of the N-1s and C-1s electrons, which is recorded on a shot-by-shot basis. Figure 2A shows the difference image between the average over all shots corresponding to a narrow range of streaking directions sorted by post-selection of the images and the average over all streaked images without any sorting and selection. An angular shift between the streaking directions of C-1s (outer) and N-1s (inner) electrons can be observed, which directly reflects the relative photoionization delays, as described above. To accurately determine the relative streaking angle, we employ a partial covariance analysis that utilizes the large shot-to-shot variation of streaking directions to directly extract without estimating and sorting the streaking directions of C-1s and N-1s electrons in every single image. For a given relative streaking angle , the streaking-induced signal fluctuations in C-1s electrons at have the most positive covariance with those in N-1s electrons around . By correctly calculating the partial covariance coefficients between streaking-induced signal changes in N-1s and C-1s electrons, the differential streaking angle manifests itself as a global shift in the partial covariance maps along the diagonal, as shown in Fig. 2C. Additional details on the partial covariance analysis are given in Section 2.2.1 of the supplementary material (SM). By comparing Figs. 3A,C to B,D, a relative photoionization delay of up to as is observed at eV, as opposed to a negligible delay at eV. We have also developed a complementary data-analysis method (the so-called ”center-of-mass” method), which is described in Section 2.2.2 of the SM. These two independent analysis methods yielded consistent photoionization delays (Fig. S14), which confirms the reliability and robustness of both methods.
Figure 3 summarizes all photoionization delays measured in this work. Panel A shows the measured photoionization delays between the N-1s and C-1s photoelectrons, which are excellent approximations to the absolute photoionization delays of the N-1s electrons. These delays are compared to the absolute N-1s photoionization delays from core-level photoionization calculations using the Schwinger variational principle with a correlation-polarization potential, summed with the Coulomb-laser coupling (CLC) delays. Details on the calculations are given in Section 3 of the SM. The delays shown in Fig. 3A globally decrease with increasing kinetic energy, which reflects the decreasing sensitivity of faster electrons to the potential of the molecular cations. The delays of all three molecules feature local maxima between 408 and 420 eV, which are in reasonable agreement with the local maxima in the calculations. These local maxima in both experiment and theory originate from shape resonances, which correspond to a transient trapping of the photoelectron for a few tens to hundreds of attoseconds before it escapes from the molecule.
Since such shape resonances are notoriously sensitive to electron correlations and molecular structure (?, ?, ?), we have carefully benchmarked the core-level photoionization calculations against measured photoionization cross sections (inset of Fig. 3A and Fig. S22), as well as photoelectron angular distributions measured in the present work (Figs. 3C and S23). Comparing the calculations using the equivalent-core (ECO) method with the experimental photoionization cross section (inset of Fig. 3A) revealed very good agreement for the resonance lying below the nitrogen K-edge and the lowest-lying shape resonance after shifting the calculations to lower energies by 4.3 eV. In comparison, the correlation-polarization (CP) calculations only needed to be shifted by 0.6 eV to agree with the experimental cross sections and the calculated width of the lowest-lying shape resonance was in better agreement with the experiment than the ECO calculations. The CP calculations also agree better with the angular asymmetry parameters measured in this work (Figs. 3C and S23). Since both calculations overestimate the energetic position of the upper shape resonance by 3.5 eV, we therefore identified the CP calculations as the more accurate ones and from hereon exclusively discuss these results.
Figure 3B compares the calculated photoionization delays (full lines), with and without the convolution used to account for the 3.8 eV bandwidth of the attosecond X-ray pulses (Fig. S1), to the photoionization cross sections (dashed lines). The two local maxima in each curve correspond to shape resonances that have been further characterized in separate calculations, detailed in Section 3.2 of the SM. The insets show the one-electron representations of the resonant wave functions (Kapur-Peierls states, see SM) corresponding to these two shape resonances. In both cases, electron trapping occurs in the molecular plane with the resonances displaying an anti-bonding character. They are therefore mirror-symmetric with respect to the molecular plane and belong to the a symmetry in the C point group of the single-site core-ionized pyrimidine. This is confirmed by the fact that both local maxima appear in the cross section corresponding to a total final-state wave function (ion+electron) of A symmetry. The lower-lying shape resonance features 5 nodal planes and the upper one 6, showing that partial waves of high angular momentum are dominating the photoelectron trapping in both cases. The narrower width of the lower resonance nicely correlates with a longer photoemission delay and the broader linewidth of the upper one correlates with the shorter delay. The partial cross section corresponding to a total final-state wavefunction of A symmetry, i.e., anti-symmetric with respect to the molecular plane and thus a -type continuum wave function, does not feature any sharp resonances, indicating that photoelectron trapping exclusively occurs in the molecular plane of pyrimidine. Similar results have been obtained for pyridine and -triazine and are shown in Fig. S20.
Since our measurements extend to within 2 eV of the N-1s ionization threshold, they are particularly sensitive to the CLC delays. It has previously been shown that CLC delays in attosecond streaking experiments are equivalent to the continuum-continuum (cc) delays known from RABBIT measurements (?). The most commonly used approximation, known as “P”, only includes the phase of the continuum-continuum (cc) matrix elements (?). The “P+A” formula additionally contains a long-range amplitude correction. Through detailed comparisons of these (and two more) formulae with calculations based on the time-dependent Schrödinger equation (TDSE, see SM, Fig. S24A), we found that the arithmetic average of “P” and “P+A” agreed best with the TDSE. We therefore used those results in Figs. 3 and 4 of the present article and show the other results in the remainder of Fig. S24. We also investigated the role of the Auger decay and the post-collision interaction (PCI) between the Auger electron and the photoelectron. We find that the correction of the Coulomb delay from a singly- to a doubly-charged cation is nearly compensated by the delay that is caused by the PCI effect (Fig. S25). Since these corrections are small and do not affect the relative delays between the molecular species that we studied, we refer the interested reader to the SM, and neglect these corrections for the remainder of the present discussion.
These results open the interesting opportunity to analyze the mechanisms that govern the trapping of the photoelectron on the attosecond time scale. Figure 4A shows a sketch of effective radial potentials which indicate the height of the angular-momentum barrier for different values of the angular momentum, , and Figs. 4C,E,G show the absolute squares of resonant one-electron wavefunctions, resolved into partial waves of angular momentum and irreducible representations of the respective point groups (specified in each panel). Figure 4B shows the CLC delays and Figs. 4D,F,H show the calculated delays for all three molecules. Because of its isolated nature and visibility in the experimental results (Fig. 3A), we focus the remainder of this discussion on the shape resonance located at 417-420 eV in each molecule.
In all three molecules, the partial wave(s) with dominate(s) in amplitude, which is the consequence of the high barrier in the radial potential (panel A) arising from the centrifugal-potential contribution , where is the mass of the electron and its distance from the center of charge. In the case of pyridine (Fig. 4A), the dominant partial wave is least-well trapped compared to the other molecules, which manifests itself in the ratio of the absolute-square amplitudes of the resonant wave function between the local maximum at small and the local minimum indicated by the second grey arrow in Figs. 4C,E,G. This identifies tunneling through an angular-momentum barrier as the first mechanism governing electron trapping in the azabenzene series and shows that the trapping strength of the dominant partial wave increases in the order pyridine, pyrimidine, -triazine.
Figure 4 moreover reveals a second mechanism at play. In addition to the changing amplitude ratio between the quasi-bound and outer parts of the wave functions, we also observe the manifestation of molecular symmetry. Whereas pyridine displays sizable contributions from an partial wave, which can be identified as being weakly trapped from the very low amplitude of its quasi-bound part, the contributions of are much weaker in pyrimidine and barely visible in -triazine as a consequence of symmetry. The three-fold (D) symmetry of the -triazine scaffold is indeed the reason why and partial waves are dominant and other partial waves are suppressed in that case. The suppression of these lower-, more weakly trapped partial waves, to which the partial waves can couple to release the photoelectron more efficiently, explains the increasing lifetime of the shape resonance in the series pyridine, pyrimidine, -triazine.
To summarize, these two mechanisms explain the trend in the resonant photoelectron trapping times in the azabenzene series. This trend is most clearly visualized in Figs. 4D,F,H, which show that the photoionization delays in the vicinity of the discussed shape resonance indeed increase in the sequence pyridine, pyrimidine, -triazine.
In conclusion, we have used the novel attosecond capabilities of LCLS to measure core-level photoionization delays of aromatic molecules. These measurements were made possible by the combination of attosecond X-ray pulses with the angular streaking technique. Absolute N-1s photoionization delays were accessed by using the much faster C-1s photoelectrons as a time reference. These self-referenced measurements were realized on a fine energy grid and revealed the presence of local maxima at 3-15 eV above the N-1s threshold. Comparison to advanced core-level photoionization calculations identified two dominant shape resonances in each molecule and showed that photoelectron trapping occurs in the plane of the aromatic molecules. The two dominant mechanisms governing the photoelectron trapping were identified as (i) changes in the barrier height of the molecular potential that lead to an increase of the electron trapping with increasing number of nitrogen atoms and (ii) the increasing symmetry of the molecular scaffold reduces the coupling between the well-trapped high- and the less-trapped lower- states. Both effects individually and collectively increase the trapping time in the sequence pyridine, pyrimidine, -triazine. These results demonstrate the promise of performing attosecond core-level photoemission experiments at free-electron-laser facilities for the following reasons. The studied molecules are indeed key building blocks of biologically active molecules, as well as molecules employed in optoelectronics, which opens the perspective of studying attosecond electron dynamics in these systems. The key advantage of core-level photoionization delays over valence-shell photoionization delays, is that the photoelectron spectra always remain simple, featuring only one dominant peak per atom of a given type. This will enable the application of the methods described in this work to complex molecules, molecular assemblies and polymers, such as organic semiconductors. Specifically, the effects governing the attosecond electron trapping in aromatic systems unveiled in the present work may become guiding principles for the design and interpretation of experiments studying charge migration and charge transfer in functional materials.
Acknowledgments
Funding: Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. The work of J.-B.J., X.G., D.H., M.H., S.H. and H.J.W. was supported by ETH Zürich and project 200021_172946 of the Swiss National Science Foundation. The work of M.H. was additionally supported by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 801459 - FP-RESOMUS. The effort from T.D.D., J.T.O. A.L.W., J.W., T.J.A.W and J.P.C. was supported by DOE, BES, Chemical Sciences, Geosciences, and Biosciences Division (CSGB). Z.G, D.C., J.D., P.L.F., R.R. and A.M. acknowledge support from the Accelerator and Detector Research Program of the Department of Energy, Basic Energy Sciences division. Z.G., P.L.F. and R.R. also acknowledge support from Robert Siemann Fellowship of Stanford University. The effort by R.R.L. and C.W.M. at LBNL was supported by the U.S. DOE BES, CSGB under Contract No. DE-AC02-05CH11231. Calculations at LBNL made use of the resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility, and the Lawrencium computational cluster resource provided by the IT Division at the LBNL. C.S.T. was supported in part by the Berkeley Lab Undergraduate Faculty Fellowship (BLUFF) Program, managed by Workforce Development & Education at Berkeley Lab.
Author Contributions: J.-B.J., T.D., K.U., A.M., J.P.C. and H.J.W. conceived the experiment. All authors contributed to the realization of the experiment and discussed the results. J.-B.J., Z.G. and T.D. developed the data-analysis methodology. J.-B.J. and Z.G. analyzed the data. R.R.L., C.S.T. and C.W.M. performed the calculations of molecular photoionization delays. J.-B.J., Z.G. and H.J.W. wrote the initial draft of the manuscript, which was reviewed and edited by all authors. Competing Interests: The authors declare no competing interests. Data availability statement: All data will be made available on the public ETH data repository upon acceptance of the manuscript.

### v3

@@
-Attosecond Core-level Photoionization Delays of Aromatic Molecules
-Attosecond Core-level Photoionization Delays
-of Aromatic Molecules
-Attosecond photoionization delays are a unique probe of the structure and the electronic dynamics of matter, but valence-shell photoionization dynamics of polyatomic systems are often complicated and challenging to interpret because of the delocalized nature of the electronic wave functions.
-Here we use attosecond soft-X-ray pulses from LCLS to probe element-specific core-level photoionization delays between N-1s and C-1s core shells of aromatic azabenzene molecules containing one to three nitrogen atoms.
-Using a self-referenced technique, we measure absolute N-1s photoionization delays of up to 300 as, which increase with the number of nitrogen atoms in the molecule.
-Comparisons to molecular photoionization calculations reveal the dominant mechanisms to be the increasing trapping of the photoelectron wavefunction in the molecular plane, as well as the symmetry-induced decrease of the coupling strength among the partial waves.
-This study demonstrates the unique opportunities opened by measurements of core-level photoionization delays for unravelling attosecond electronic dynamics in the building blocks of biomolecules and molecular optoelectronics.
-Photoionization or photoemission is often conceptualized as a half-scattering event between a photoelectron wave and matter (?, ?).
-Owing to the atomic length scale of the electron de-Broglie wave, it is a sensitive probe of the atomic structure of matter (?).
-Recently, attosecond metrology has added the unique aspect of temporal resolution on atomic time scales, which is absent from traditional collision physics.
-This advance (?, ?), which has been highlighted in the 2023 Nobel Prize in Physics (?), has integrated time-domain access to multi-electron dynamics into the powerful framework of electron-scattering physics.
-Not unexpectedly, these unique opportunities have created a flourishing research field with applications to molecules (?, ?, ?, ?, ?, ?), clusters (?, ?), liquids (?) and solids (?, ?, ?, ?).
-One conclusion from these recent studies is that photoionization delays are sensitive to the spatial extension of electronic wave functions (?).
-Whereas this property offers exciting avenues for studying the dynamics of electron-hole (de)localization dynamics (?), it may obscure other mechanisms governing photoionization delays because of the strong spatial delocalization characteristic of the valence shell.
-Moreover, the delocalized nature of valence-shell wavefunctions prevents atomic-scale resolution of the photoionization event and the spectral congestion limits the complexity of the systems that can be studied (?).
-Here, we demonstrate the unique opportunities that arise by transposing the research field of attosecond chronoscopy (?) from the extreme-ultraviolet to the X-ray domain.
+Attosecond X-ray Chronoscopy of Core-level Photoemission
+Attosecond X-ray Chronoscopy of Core-level Photoemission
+Attosecond photoemission or photoionization delays are a unique probe of the structure and the electronic dynamics of matter.
+However, spectral congestion and spatial delocalization of valence electron wave functions set fundamental limits to the complexity of systems that can be studied and the information that can be retrieved, respectively.
+Using attosecond X-ray pulses from LCLS, we demonstrate the key advantages of measuring core-level delays: the photoelectron spectra remain atom-like, the measurements become element specific and the observed scattering dynamics originate from a point-like source.
+We exploit these unique features to reveal the effects of electronegativity and symmetry on attosecond scattering dynamics by measuring the photoionization delays between N-1s and C-1s core shells of a series of aromatic azabenzene molecules.
+Remarkably, the delays systematically increase with the number of nitrogen atoms in the molecule and reveal multiple resonances.
+We identify two previously unknown mechanisms regulating the associated attosecond dynamics, namely the enhanced confinement of the trapped wavefunction with increasing electronegativity of the atoms and the decrease of the coupling strength among the photoemitted partial waves with increasing symmetry.
+This study demonstrates the unique opportunities opened by measurements of core-level photoionization delays for unravelling attosecond electron dynamics in complex matter.
+Tracking attosecond multi-electron dynamics with atomically resolved spatial information is the current focus of numerous research efforts worldwide.
+X-ray spectroscopy has established itself as a powerful solution to breaking the complexity barriers faced by valence-shell spectroscopies.
+These barriers are particularly noticeable in time-resolved studies of the photoelectric effect, known as attosecond chronoscopy (?, ?, ?).
+This technique has integrated time-domain access to multi-electron dynamics into the powerful framework of electron-scattering physics which has created a flourishing research field with applications to molecules (?, ?, ?, ?, ?, ?), clusters (?, ?), liquids (?) and solids (?, ?, ?, ?).
+Such studies have traditionally made use of attosecond pulses in the extreme ultraviolet (up to 120 eV), which has so far prevented access to atomic core levels.
+Here, we demonstrate the unique opportunities that arise from advancing the research field of attosecond chronoscopy from the extreme-ultraviolet to the X-ray domain.
 This adds element specificity and site selectivity to the method, while simultaneously confining the emission site of the photoelectron wave to a point-like source.
-Moreover, core-shell photoelectron spectra have a much simpler structure than valence-shell spectra because they feature a single band per atom of each type (when neglecting satellite lines), which facilitates the extension of the method to complex samples.
-This advance is made possible by groundbreaking progress in free-electron laser (FEL) science that has led to the generation of attosecond pulses in the X-ray domain through the XLEAP technique (?).
-The long-standing challenge of temporal jitter between the FEL pulses and laser pulses is solved by using attosecond angular streaking with a circularly-polarized infrared laser pulse and detecting the photoelectron momentum distribution in the polarization plane (?, ?).
-The present experiments aim at a fundamental understanding of the attosecond dynamics of electron trapping in aromatic systems (i.e., benzene derivatives), specifically the multiple scattering phenomena that lead to transient trapping of photoelectrons following their release from specific atoms.
-Our results provide answers to important questions on mechanisms that regulate electron trapping, and the duration over which electrons are trapped.
-For this research, we chose the azabenzene molecules pyridine, pyrimidine and -triazine, depicted in Fig.
+Moreover, core-shell photoelectron spectra have a much simpler structure than valence-shell spectra because they feature a single band per atom of each type (when neglecting satellite lines), which facilitates the extension of the method to complex forms of matter.
+This advance was made possible by groundbreaking progress in free-electron-laser (FEL) science that has led to the generation of attosecond pulses in the X-ray domain through the XLEAP technique (?).
+The long-standing challenge of temporal jitter between the FEL pulses and laser pulses was solved by using attosecond angular streaking with a circularly-polarized infrared laser pulse and detecting the photoelectron momentum distribution in the polarization plane (?, ?).
+The present experiments establish the unique capabilities of X-ray chronoscopy at unraveling the mechanisms of attosecond electron dynamics in complex systems, notably in aromatic molecules (i.e., benzene derivatives).
+Specifically, our study reveals the multiple-scattering phenomena that lead to transient trapping of photoelectrons following their release from specific atoms.
+Our results reveal two previously unknown mechanisms regulating such attosecond scattering dynamics.
+First, we find that substituting the ionized functional group with a more electronegative one (C-H N) increases the trapping times.
+Second, we show that increasing the symmetry of the molecular scaffold also increases the trapping times by suppressing the coupling to the less-trapped partial waves with lower angular momentum.
+To demonstrate these effects, we chose the azabenzene molecules pyridine, pyrimidine and -triazine, depicted in Fig.
 1A.
@@
 Since the latter have kinetic energies of more than 110 eV, the C-1s electrons are emitted with negligible photoionization delays ( as), defining a self-referenced technique for accessing absolute photoionization delays of the N-1s electrons.
-We measure N-1s photoionization delays of up to 300 as close to the nitrogen -shell threshold, which generally decrease with increasing kinetic energy but display local maxima at 3-15 eV above the threshold.
+We measured N-1s photoionization delays of up to 300 as close to the nitrogen -shell threshold, which generally decrease with increasing kinetic energy but display local maxima at 3-15 eV above the threshold.
 The measured delays also markedly increase from pyridine to pyrimidine and -triazine.
@@
 Additionally, the increasing symmetry of the molecular scaffold (from C in pyridine to D in -triazine) reduces the coupling among the photoelectron partial waves, which frustrates the decay of the trapped partial waves with high angular momentum.
-Our study reveals two important mechanisms regulating electronic dynamics in aromatic systems, i.e., the dependence of photoelectron trapping on molecular structure and molecular symmetry.
 Pyrimidine is a building block of the nucleobases thymine, cytosine and uracil.
@@
 Figure 3 summarizes all photoionization delays measured in this work.
-Panel A shows the measured photoionization delays between the N-1s and C-1s photoelectrons, which are excellent approximations to the absolute photoionization delays of the N-1s electrons.
+Panels A, C, and E show the measured photoionization delays between the N-1s and C-1s photoelectrons for the three molecules, which are excellent approximations to the absolute photoionization delays of the N-1s electrons.
 These delays are compared to the absolute N-1s photoionization delays from core-level photoionization calculations using the Schwinger variational principle with a correlation-polarization potential, summed with the Coulomb-laser coupling (CLC) delays.
 Details on the calculations are given in Section 3 of the SM.
-The delays shown in Fig.
-3A globally decrease with increasing kinetic energy, which reflects the decreasing sensitivity of faster electrons to the potential of the molecular cations.
+The comparison of the experimental and theoretical delays among the three molecules can be found in panel F and the inset of panel A, respectively.
+The delays globally decrease with increasing kinetic energy, which reflects the decreasing sensitivity of faster electrons to the potential of the molecular cations.
 The delays of all three molecules feature local maxima between 408 and 420 eV, which are in reasonable agreement with the local maxima in the calculations.
 These local maxima in both experiment and theory originate from shape resonances, which correspond to a transient trapping of the photoelectron for a few tens to hundreds of attoseconds before it escapes from the molecule.
-Since such shape resonances are notoriously sensitive to electron correlations and molecular structure (?, ?, ?), we have carefully benchmarked the core-level photoionization calculations against measured photoionization cross sections (inset of Fig.
-3A and Fig.
-S22), as well as photoelectron angular distributions measured in the present work (Figs.
-3C and S23).
-Comparing the calculations using the equivalent-core (ECO) method with the experimental photoionization cross section (inset of Fig.
-3A) revealed very good agreement for the resonance lying below the nitrogen K-edge and the lowest-lying shape resonance after shifting the calculations to lower energies by 4.3 eV.
+Since such shape resonances are notoriously sensitive to electron correlations and molecular structure (?, ?, ?), we have carefully benchmarked the core-level photoionization calculations against measured photoionization cross sections (Figs.
+3D and S22), as well as photoelectron angular distributions measured in the present work (Figs.
+3B and S23).
+Comparing the calculations using the equivalent-core (ECO) method with the experimental photoionization cross section (Fig.
+3D) revealed very good agreement for the resonance lying below the nitrogen K-edge and the lowest-lying shape resonance after shifting the calculations to lower energies by 4.3 eV.
 In comparison, the correlation-polarization (CP) calculations only needed to be shifted by 0.6 eV to agree with the experimental cross sections and the calculated width of the lowest-lying shape resonance was in better agreement with the experiment than the ECO calculations.
 The CP calculations also agree better with the angular asymmetry parameters measured in this work (Figs.
-3C and S23).
+3B and S23).
 Since both calculations overestimate the energetic position of the upper shape resonance by 3.5 eV, we therefore identified the CP calculations as the more accurate ones and from hereon exclusively discuss these results.
-Figure 3B compares the calculated photoionization delays (full lines), with and without the convolution used to account for the 3.8 eV bandwidth of the attosecond X-ray pulses (Fig.
-S1), to the photoionization cross sections (dashed lines).
-The two local maxima in each curve correspond to shape resonances that have been further characterized in separate calculations, detailed in Section 3.2 of the SM.
-The insets show the one-electron representations of the resonant wave functions (Kapur-Peierls states, see SM) corresponding to these two shape resonances.
-In both cases, electron trapping occurs in the molecular plane with the resonances displaying an anti-bonding character.
-They are therefore mirror-symmetric with respect to the molecular plane and belong to the a symmetry in the C point group of the single-site core-ionized pyrimidine.
-This is confirmed by the fact that both local maxima appear in the cross section corresponding to a total final-state wave function (ion+electron) of A symmetry.
+Figures 3C,D demonstrate the two shape resonances of pyrimidine within the energy range, which lead to the local maxima of the cross section.
+The A symmetry component is dominating in the cross section, which displays an anti-bonding character and corresponds to the electron trapping that occurs in the molecular plane in both resonances.
+The insets in Fig.
+3C show the one-electron representations of the corresponding resonant wavefunctions (Kapur-Peierls states, see SM).
 The lower-lying shape resonance features 5 nodal planes and the upper one 6, showing that partial waves of high angular momentum are dominating the photoelectron trapping in both cases.
@@
 The partial cross section corresponding to a total final-state wavefunction of A symmetry, i.e., anti-symmetric with respect to the molecular plane and thus a -type continuum wavefunction, does not feature any sharp resonances, indicating that photoelectron trapping exclusively occurs in the molecular plane of pyrimidine.
-Similar results have been obtained for pyridine and -triazine and are shown in Fig.
+The probability density of the resonant wavefunction of -triazine at 418 eV is illustrated in the inset of Fig.
+3E and in Fig.
+4.
+Similar resonances occur in all three molecules, as shown in Fig.
 S20.
@@
 The “P+A” formula additionally contains a long-range amplitude correction.
-Through detailed comparisons of these (and two more) formulae with calculations based on the time-dependent Schrödinger equation (TDSE, see SM, Fig.
+Through detailed comparisons of these (and two more) formulae with calculations based on the time-dependent Schrödinger equation (TDSE, see Fig.
 S24A), we found that the arithmetic average of “P” and “P+A” agreed best with the TDSE.
@@
 Since these corrections are small and do not affect the relative delays between the molecular species that we studied, we refer the interested reader to the SM, and neglect these corrections for the remainder of the present discussion.
-These results open the interesting opportunity to analyze the mechanisms that govern the trapping of the photoelectron on the attosecond time scale.
-Figure 4A shows a sketch of effective radial potentials which indicate the height of the angular-momentum barrier for different values of the angular momentum, , and Figs.
-4C,E,G show the absolute squares of resonant one-electron wavefunctions, resolved into partial waves of angular momentum and irreducible representations of the respective point groups (specified in each panel).
+These results open the novel opportunity to analyze the mechanisms that govern the trapping of the photoelectron on the attosecond time scale.
+Figure 4 illustrates the trapping of the outgoing photoelectron wave in the higher-lying (418 eV) resonance of -triazine.
+The atomic cores provide attractive potentials in their vicinities, while the centrifugal energy creates a repulsive region at the center of the aromatic ring.
+Their addition results in a potential barrier (black curve in Fig.
+4) which traps the resonant wavefunction for a duration of 360 as.
+The photoelectron wave initially originates from the 1s shells of the nitrogen atoms and then scatters in the molecular potential, which results in comparable amplitudes of the resonant wave function on all atoms of the molecular ring.
+But what causes the increase of the photoionization delays and resonance lifetimes with increasing number of nitrogen atoms and increasing symmetry?
+To answer this question, Figure 5A shows a sketch of effective radial potentials which indicate the height of the angular-momentum barrier for different values of the angular momentum, , and Figs.
+5C,E,G show the absolute squares of resonant one-electron wavefunctions, resolved into partial waves of angular momentum and irreducible representations of the respective point groups (specified in each panel).
 Figure 5B shows the CLC delays and Figs.
-4D,F,H show the calculated delays for all three molecules.
+5D,F,H show the calculated delays for all three molecules.
 Because of its isolated nature and visibility in the experimental results (Fig.
-3A), we focus the remainder of this discussion on the shape resonance located at 417-420 eV in each molecule.
+3C), we focus the remainder of this discussion on the shape resonance located at 417-420 eV in each molecule.
 In all three molecules, the partial wave(s) with dominate(s) in amplitude, which is the consequence of the high barrier in the radial potential (panel A) arising from the centrifugal-potential contribution , where is the mass of the electron and its distance from the center of charge.
+The effective potentials of are indicated by the shaded areas in the corresponding panels.
+It is clear that the main part of the wavefunction is confined in the quasi-bound region in the vicinity of the N and C atoms, while a tail tunnels through the barrier.
 In the case of pyridine (Fig.
-4A), the dominant partial wave is least-well trapped compared to the other molecules, which manifests itself in the ratio of the absolute-square amplitudes of the resonant wave function between the local maximum at small and the local minimum indicated by the second grey arrow in Figs.
-4C,E,G.
+5A), the dominant partial wave is least-well trapped compared to the other molecules, which manifests itself in the ratio of the absolute-square amplitudes of the resonant wavefunction between the local maximum at small and the local minimum indicated by the second grey arrow in Figs.
+5C,E,G.
 This identifies tunneling through an angular-momentum barrier as the first mechanism governing electron trapping in the azabenzene series and shows that the trapping strength of the dominant partial wave increases in the order pyridine, pyrimidine, -triazine.
-Figure 4 moreover reveals a second mechanism at play.
-In addition to the changing amplitude ratio between the quasi-bound and outer parts of the wave functions, we also observe the manifestation of molecular symmetry.
+This shows that substituting a C-H group with the more electronegative N atom increases the trapping strength and therefore the lifetime of the resonance.
+We note that the divergence of the Kapur-Peierls state for large is a consequence of their mathematical nature, specifically their complex-valued momentum, see SM Section 3.2.
+Figure 5 also reveals a second mechanism at play.
+In addition to the changing amplitude ratio between the wavefunctions in the quasi-bound and tunneling regions, we also observe the manifestation of molecular symmetry.
 Whereas pyridine displays sizable contributions from an partial wave, which can be identified as being weakly trapped from the very low amplitude of its quasi-bound part, the contributions of are much weaker in pyrimidine and barely visible in -triazine as a consequence of symmetry.
@@
 This trend is most clearly visualized in Figs.
-4D,F,H, which show that the photoionization delays in the vicinity of the discussed shape resonance indeed increase in the sequence pyridine, pyrimidine, -triazine.
+5D,F,H, which show that the photoionization delays in the vicinity of the discussed shape resonance indeed increase in the sequence pyridine, pyrimidine, -triazine.
 In conclusion, we have used the novel attosecond capabilities of LCLS to measure core-level photoionization delays of aromatic molecules.
@@
 Comparison to advanced core-level photoionization calculations identified two dominant shape resonances in each molecule and showed that photoelectron trapping occurs in the plane of the aromatic molecules.
-The two dominant mechanisms governing the photoelectron trapping were identified as (i) changes in the barrier height of the molecular potential that lead to an increase of the electron trapping with increasing number of nitrogen atoms and (ii) the increasing symmetry of the molecular scaffold reduces the coupling between the well-trapped high- and the less-trapped lower- states.
+The two dominant mechanisms governing the photoelectron trapping were identified as (i) changes in the barrier height of the molecular potential that lead to an increase of the electron trapping with increasing electronegativity of the functional groups and (ii) the increasing symmetry of the molecular scaffold reduces the coupling between the well-trapped high- and the less-trapped lower- states.
 Both effects individually and collectively increase the trapping time in the sequence pyridine, pyrimidine, -triazine.

### v4

@@
-Attosecond X-ray Chronoscopy of Core-level Photoemission
-Attosecond X-ray Chronoscopy of Core-level Photoemission
+Attosecond X-ray Core-level Chronoscopy of Aromatic Molecules
+Abstract
 Attosecond photoemission or photoionization delays are a unique probe of the structure and the electronic dynamics of matter.
@@
 Using attosecond X-ray pulses from LCLS, we demonstrate the key advantages of measuring core-level delays: the photoelectron spectra remain atom-like, the measurements become element specific and the observed scattering dynamics originate from a point-like source.
-We exploit these unique features to reveal the effects of electronegativity and symmetry on attosecond scattering dynamics by measuring the photoionization delays between N-1s and C-1s core shells of a series of aromatic azabenzene molecules.
-Remarkably, the delays systematically increase with the number of nitrogen atoms in the molecule and reveal multiple resonances.
+We exploit these unique features to reveal the effects of electronegativity and symmetry on attosecond scattering dynamics by measuring and calculating the photoionization delays between N-1s and C-1s core shells of a series of aromatic azabenzene molecules.
+Remarkably, the delays increase with the number of nitrogen atoms in the molecule and reveal multiple resonances.
 We identify two previously unknown mechanisms regulating the associated attosecond dynamics, namely the enhanced confinement of the trapped wavefunction with increasing electronegativity of the atoms and the decrease of the coupling strength among the photoemitted partial waves with increasing symmetry.
-This study demonstrates the unique opportunities opened by measurements of core-level photoionization delays for unravelling attosecond electron dynamics in complex matter.
+This study demonstrates the unique opportunities opened by measurements of core-level photoionization delays for unraveling attosecond electron dynamics in complex matter.
+I Introduction
 Tracking attosecond multi-electron dynamics with atomically resolved spatial information is the current focus of numerous research efforts worldwide.
 X-ray spectroscopy has established itself as a powerful solution to breaking the complexity barriers faced by valence-shell spectroscopies.
-These barriers are particularly noticeable in time-resolved studies of the photoelectric effect, known as attosecond chronoscopy (?, ?, ?).
-This technique has integrated time-domain access to multi-electron dynamics into the powerful framework of electron-scattering physics which has created a flourishing research field with applications to molecules (?, ?, ?, ?, ?, ?), clusters (?, ?), liquids (?) and solids (?, ?, ?, ?).
+These barriers are particularly pronounced in time-resolved studies of the photoelectric effect, known as attosecond chronoscopy [1, 2, 3].
+This technique has integrated time-domain access to multi-electron dynamics into the powerful framework of electron-scattering physics which has created a flourishing research field with applications to molecules [4, 5, 6, 7, 8, 9], clusters [10, 11], liquids [12] and solids [13, 14, 15, 16].
 Such studies have traditionally made use of attosecond pulses in the extreme ultraviolet (up to 120 eV), which has so far prevented access to atomic core levels.
@@
 This adds element specificity and site selectivity to the method, while simultaneously confining the emission site of the photoelectron wave to a point-like source.
-Moreover, core-shell photoelectron spectra have a much simpler structure than valence-shell spectra because they feature a single band per atom of each type (when neglecting satellite lines), which facilitates the extension of the method to complex forms of matter.
-This advance was made possible by groundbreaking progress in free-electron-laser (FEL) science that has led to the generation of attosecond pulses in the X-ray domain through the XLEAP technique (?).
-The long-standing challenge of temporal jitter between the FEL pulses and laser pulses was solved by using attosecond angular streaking with a circularly-polarized infrared laser pulse and detecting the photoelectron momentum distribution in the polarization plane (?, ?).
-The present experiments establish the unique capabilities of X-ray chronoscopy at unraveling the mechanisms of attosecond electron dynamics in complex systems, notably in aromatic molecules (i.e., benzene derivatives).
+Moreover, core-shell photoelectron spectra have a much simpler structure than valence-shell spectra, which facilitates the extension of the method to complex forms of matter.
+This advance was made possible by groundbreaking progress in free-electron-laser (FEL) science that has led to the generation of attosecond pulses in the X-ray domain through the X-ray laser-enhanced attosecond pulse generation (XLEAP) technique [17].
+The long-standing challenge of temporal jitter between the FEL pulses and laser pulses was solved by using attosecond angular streaking with a circularly-polarized infrared laser pulse and detecting the photoelectron momentum distribution in the polarization plane [18, 19, 20].
+The present experiments establish the unique capabilities of X-ray chronoscopy, recently demonstrated in a diatomic molecule [20], at unraveling the mechanisms of attosecond electron dynamics in complex systems, notably in aromatic molecules (i.e., benzene derivatives).
 Specifically, our study reveals the multiple-scattering phenomena that lead to transient trapping of photoelectrons following their release from specific atoms.
@@
 1A.
-We produced attosecond X-ray pulses with energies just above the nitrogen K-edge located at 405 eV and measured the photoionization delays between the electrons emitted from the nitrogen -shell (N-1s) and those emitted from the carbon -shell (C-1s).
+We produced attosecond X-ray pulses with energies just above the nitrogen K-edge located at 405 eV and measured the photoionization delays between the electrons emitted from the nitrogen K-shell (N-1s) and those emitted from the carbon K-shell (C-1s).
 Since the latter have kinetic energies of more than 110 eV, the C-1s electrons are emitted with negligible photoionization delays ( as), defining a self-referenced technique for accessing absolute photoionization delays of the N-1s electrons.
-We measured N-1s photoionization delays of up to 300 as close to the nitrogen -shell threshold, which generally decrease with increasing kinetic energy but display local maxima at 3-15 eV above the threshold.
-The measured delays also markedly increase from pyridine to pyrimidine and -triazine.
+We measured N-1s photoionization delays of up to 300 as close to the nitrogen K-shell threshold, which generally decrease with increasing kinetic energy but display local maxima at 3-15 eV above the threshold.
+The measured delays indeed markedly increase from pyridine to pyrimidine.
 These experimental results are compared with state-of-the-art core-level photoionization calculations, which reveal the presence of several shape resonances in this low kinetic-energy region.
-The comparison of experiment and theory reveals that the photoelectrons originating from the nitrogen -shell are trapped with increasing efficiency in the molecular plane in the sequence pyridine, pyrimidine, -triazine, i.e., with increasing number of nitrogen atoms.
-Additionally, the increasing symmetry of the molecular scaffold (from C in pyridine to D in -triazine) reduces the coupling among the photoelectron partial waves, which frustrates the decay of the trapped partial waves with high angular momentum.
+The comparison of experiment and theory reveals that the photoelectrons originating from the nitrogen K-shell are trapped with increasing efficiency in the molecular plane in the sequence pyridine, pyrimidine, -triazine, i.e., with increasing number of nitrogen atoms.
+Additionally, the increasing symmetry of the molecular scaffold (from C2v in pyridine to D3h in -triazine) reduces the coupling among the photoelectron partial waves, which frustrates the decay of the trapped partial waves with high angular momentum.
 Pyrimidine is a building block of the nucleobases thymine, cytosine and uracil.
-All three azabenzene molecules are widespread structural motifs in biomolecules, drugs and molecular optoelectronics (?, ?, ?, ?).
-The discovered mechanisms therefore suggest guidelines for understanding electronic dynamics in biomolecules and molecular optoelectronics.
-Figure 1 illustrates the principle of the present experiments, carried out at the time-resolved atomic, molecular, and optical science (TMO) experimental hutch (?) of the Linac Coherent Light Source (LCLS).
-Linearly polarized attosecond X-ray pulses with a tunable central energy were obtained using the XLEAP technique (?) and were superimposed with circularly polarized short-wave infrared (IR, 1.3 m, 40 fs) pulses in the interaction region of a co-axial velocity-map-imaging (c-VMI) spectrometer.
+All three azabenzene molecules are widespread structural motifs in biomolecules, drugs and molecular optoelectronics [21, 22, 23, 24].
+The discovered mechanisms therefore suggest guidelines for extending attosecond chronoscopy to biomolecules and molecules used in optoelectronics.
+II Experimental setup and data analysis
+Figure 1 illustrates the principle of the present experiments, carried out at the time-resolved atomic, molecular, and optical science (TMO) experimental hutch [25] of the Linac Coherent Light Source (LCLS).
+Linearly polarized attosecond X-ray pulses with a tunable central energy were obtained using the XLEAP technique [17] and were superimposed with circularly polarized short-wave infrared (IR, 1.3 m, 40 fs) pulses in the interaction region of a co-axial velocity-map-imaging (c-VMI) spectrometer.
 A continuous molecular beam containing pyridine, pyrimidine or -triazine intersected with the overlapping X-ray and IR pulses in the interaction region.
-The photoelectrons created from nitrogen -shell ionization of the sample were streaked by the IR pulses and projected onto a position-sensitive imaging detector with a central hole.
+The photoelectrons created from nitrogen K-shell ionization of the sample were streaked by the IR pulses and projected onto a position-sensitive imaging detector with a central hole.
 A photoelectron image recorded by ionizing pyrimidine with X-ray pulses centered at 454 eV is shown in Fig.
-1A, whereby the inner (outer) ring corresponds to ionization of the nitrogen (carbon) -shell electrons.
-Figures 1B, 1C, and 1D show how the relative photoionization delays between the N-1s and C-1s photoelectrons were determined.
-Because each photoelectron was angularly streaked into the direction of the IR-laser vector potential at the instant of photoionization (inset of Fig.1A), the relative streaking angle between N-1s and C-1s electrons directly encodes their relative photoionization delay.
-The conversion from relative streaking angle to the relative photoionization delay is simply given by , where is the period of the streaking laser (?, ?, ?, ?, ?, ?, ?).
+1A, whereby the inner (outer) ring corresponds to ionization of the nitrogen (carbon) K-shell electrons.
+Figure 1C shows how the relative photoionization delays between the N-1s and C-1s photoelectrons were determined.
+Because each photoelectron was angularly streaked into the direction of the IR-laser vector potential at the instant of photoionization (Fig.1B), the relative streaking angle between N-1s and C-1s electrons directly encodes their relative photoionization delay.
+The conversion from relative streaking angle to the relative photoionization delay is simply given by , where is the period of the streaking laser [26, 27, 28, 29, 30, 31, 32].
 Figure 2 illustrates how the relative streaking angles were obtained from the experimental data.
-Due to the large temporal jitter (500 fs) between the arrival time of XFEL and IR laser pulses (?), the streaking directions of the core-shell photoelectrons were completely random from shot to shot.
+Due to the large temporal jitter (500 fs) between the arrival time of XFEL and IR laser pulses [33], the streaking directions of the core-shell photoelectrons were completely random from shot to shot.
 The information of relative photoionization time delay is therefore only encoded in the relative streaking angle of the N-1s and C-1s electrons, which is recorded on a shot-by-shot basis.
@@
 For a given relative streaking angle , the streaking-induced signal fluctuations in C-1s electrons at have the most positive covariance with those in N-1s electrons around .
-By correctly calculating the partial covariance coefficients between streaking-induced signal changes in N-1s and C-1s electrons, the differential streaking angle manifests itself as a global shift in the partial covariance maps along the diagonal, as shown in Fig.
+By correctly calculating the partial covariance coefficients between streaking-induced signal changes in N-1s and C-1s electrons, the differential streaking angle manifests itself as a global shift in the partial covariance maps, as shown in Fig.
 2C.
-Additional details on the partial covariance analysis are given in Section 2.2.1 of the supplementary material (SM).
+Additional details on the partial covariance analysis are given in Section 2.2.1 of the supplemental material (SM) and systematically explained in Ref.
+[34].
 By comparing Figs.
 3A,C to B,D, a relative photoionization delay of up to as is observed at eV, as opposed to a negligible delay at eV.
-We have also developed a complementary data-analysis method (the so-called ”center-of-mass” method), which is described in Section 2.2.2 of the SM.
+We have also developed a complementary data-analysis method (the so-called “center-of-mass” method), which is described in Section 2.2.2 of the SM.
 These two independent analysis methods yielded consistent photoionization delays (Fig.
@@
 These local maxima in both experiment and theory originate from shape resonances, which correspond to a transient trapping of the photoelectron for a few tens to hundreds of attoseconds before it escapes from the molecule.
-Since such shape resonances are notoriously sensitive to electron correlations and molecular structure (?, ?, ?), we have carefully benchmarked the core-level photoionization calculations against measured photoionization cross sections (Figs.
+III Theoretical calculations
+Since such shape resonances are notoriously sensitive to electron correlations and molecular structure [36, 4, 9, 8], we have carefully benchmarked the core-level photoionization calculations against measured photoionization cross sections (Figs.
 3D and S22), as well as photoelectron angular distributions measured in the present work (Figs.
@@
 Figures 3C,D demonstrate the two shape resonances of pyrimidine within the energy range, which lead to the local maxima of the cross section.
-The A symmetry component is dominating in the cross section, which displays an anti-bonding character and corresponds to the electron trapping that occurs in the molecular plane in both resonances.
+The 1A′ symmetry component is dominating in the cross section, which displays an anti-bonding character and corresponds to the electron trapping that occurs in the molecular plane in both resonances.
 The insets in Fig.
-3C show the one-electron representations of the corresponding resonant wavefunctions (Kapur-Peierls states, see SM).
+3C show the one-electron representations of the corresponding resonant wavefunctions (Kapur-Peierls states, see SM for definition).
 The lower-lying shape resonance features 5 nodal planes and the upper one 6, showing that partial waves of high angular momentum are dominating the photoelectron trapping in both cases.
 The narrower width of the lower resonance nicely correlates with a longer photoemission delay and the broader linewidth of the upper one correlates with the shorter delay.
-The partial cross section corresponding to a total final-state wavefunction of A symmetry, i.e., anti-symmetric with respect to the molecular plane and thus a -type continuum wavefunction, does not feature any sharp resonances, indicating that photoelectron trapping exclusively occurs in the molecular plane of pyrimidine.
+The partial cross section corresponding to a total final-state wavefunction of 1A′′ symmetry, i.e., anti-symmetric with respect to the molecular plane and thus a -type continuum wavefunction, does not feature any sharp shape resonances above the ionization threshold (404 eV) in the CP calculation.
+Note however, the shifted ECO calculation does have a narrow feature slightly below the threshold, as shown in Fig.
+3D, which is in agreement with the autoionization feature seen in a previous experimental study [35].
+Thus we conclude that in the continuum above the nitrogrn K-edge the photoelectron trapping is dominated by the -type resonances in the molecular plane of pyrimidine.
 The probability density of the resonant wavefunction of -triazine at 418 eV is illustrated in the inset of Fig.
@@
 Since our measurements extend to within 2 eV of the N-1s ionization threshold, they are particularly sensitive to the CLC delays.
-It has previously been shown that CLC delays in attosecond streaking experiments are equivalent to the continuum-continuum (cc) delays known from RABBIT measurements (?).
-The most commonly used approximation, known as “P”, only includes the phase of the continuum-continuum (cc) matrix elements (?).
-The “P+A” formula additionally contains a long-range amplitude correction.
-Through detailed comparisons of these (and two more) formulae with calculations based on the time-dependent Schrödinger equation (TDSE, see Fig.
-S24A), we found that the arithmetic average of “P” and “P+A” agreed best with the TDSE.
-We therefore used those results in Figs.
-3 and 4 of the present article and show the other results in the remainder of Fig.
-S24.
-We also investigated the role of the Auger decay and the post-collision interaction (PCI) between the Auger electron and the photoelectron.
+It has previously been shown that CLC delays in attosecond streaking experiments have the same origin of the continuum-continuum (CC) delays known from reconstruction of attosecond beating by interference of two-photon transitions (RABBIT) measurements [3].
+The most commonly used expressions for the CC delays are reported in Ref.
+[37], which is based on the hydrogen-like atoms, and the radial wavefunctions are approximated considering the asymptotic phase shift (denoted as “P”) or with additional consideration of the long-range amplitude correction (denoted as “P+A”).
+The approximation of the hydrogenic radial wavefunction causes errors compared to the time-dependent Schrödinger equation (TDSE), particularly in the low-kinetic-energy region [37], which is relevant in our work.
+Recently, Ji et al.
+[38] reported using the analytical formula with hydrogenic radial wavefunctions (confluent hypergeometric functions) [39] instead of their asymptotic forms to calculate the CC transition of the hydrogen-like atoms, which shows excellent agreement with the TDSE in the whole kinetic-energy range.
+Here, we have adapted the same approach for calculating the CLC delay with a slight modification, where for the CC transition the increase or decrease of the kinetic energy is a continuous quantity instead of being equal to the energy of one IR photon, and the time delay is defined based on its derivative.
+The details can be found in Section 4.2 of the SM.
+This new approach yields very good agreement to the TDSE values reported in Ref.
+[31] (see Fig.
+S24A).
+Therefore, we used those results in Figs.
+3 and 5.
+We also investigated the role of the Auger-Meitner decay and the post-collision interaction (PCI) between the Auger-Meitner electron and the photoelectron.
 We find that the correction of the Coulomb delay from a singly- to a doubly-charged cation is nearly compensated by the delay that is caused by the PCI effect (Fig.
@@
 Since these corrections are small and do not affect the relative delays between the molecular species that we studied, we refer the interested reader to the SM, and neglect these corrections for the remainder of the present discussion.
+IV Discussion
 These results open the novel opportunity to analyze the mechanisms that govern the trapping of the photoelectron on the attosecond time scale.
-Figure 4 illustrates the trapping of the outgoing photoelectron wave in the higher-lying (418 eV) resonance of -triazine.
-The atomic cores provide attractive potentials in their vicinities, while the centrifugal energy creates a repulsive region at the center of the aromatic ring.
-Their addition results in a potential barrier (black curve in Fig.
-4) which traps the resonant wavefunction for a duration of 360 as.
+For this purpose, we analyze the properties of the trapped photoelectron wave function and the potential that traps it.
+Figure 4 compares the trapping of the outgoing photoelectron wave in the resonances at around 410 eV of pyridine and -triazine.
+The atomic cores provide attractive potentials in their vicinities, while the centrifugal energy creates a repulsive region centered on the aromatic ring.
+Their addition results in a potential barrier (black curves in Fig.
+4) which traps the resonant wavefunction.
+The calculations provide a lifetime of 150 as for pyridine and 360 as for -triazine at this shape resonance.
+The stark difference in lifetime is reflected in the time-delay difference in the same energy region.
 The photoelectron wave initially originates from the 1s shells of the nitrogen atoms and then scatters in the molecular potential, which results in comparable amplitudes of the resonant wave function on all atoms of the molecular ring.
@@
 5D,F,H show the calculated delays for all three molecules.
-Because of its isolated nature and visibility in the experimental results (Fig.
-3C), we focus the remainder of this discussion on the shape resonance located at 417-420 eV in each molecule.
-In all three molecules, the partial wave(s) with dominate(s) in amplitude, which is the consequence of the high barrier in the radial potential (panel A) arising from the centrifugal-potential contribution , where is the mass of the electron and its distance from the center of charge.
-The effective potentials of are indicated by the shaded areas in the corresponding panels.
+We will focus the remainder of this discussion on the shape resonance located at about 410 eV in each molecule.
+In all three molecules, the partial-wave components with dominate in probability density, which is the consequence of the high barrier in the radial potential (panel A) arising from the centrifugal-potential contribution , where is the mass of the electron and its distance from the center of charge.
+The effective potentials of different partial waves are indicated by the shaded areas with different colors in the corresponding panels.
 It is clear that the main part of the wavefunction is confined in the quasi-bound region in the vicinity of the N and C atoms, while a tail tunnels through the barrier.
 In the case of pyridine (Fig.
-5A), the dominant partial wave is least-well trapped compared to the other molecules, which manifests itself in the ratio of the absolute-square amplitudes of the resonant wavefunction between the local maximum at small and the local minimum indicated by the second grey arrow in Figs.
-5C,E,G.
+5A), the dominant partial wave is least-well trapped compared to the other molecules, which is clear from comparing the modulations of the probability density at the barriers.
 This identifies tunneling through an angular-momentum barrier as the first mechanism governing electron trapping in the azabenzene series and shows that the trapping strength of the dominant partial wave increases in the order pyridine, pyrimidine, -triazine.
@@
 Figure 5 also reveals a second mechanism at play.
-In addition to the changing amplitude ratio between the wavefunctions in the quasi-bound and tunneling regions, we also observe the manifestation of molecular symmetry.
-Whereas pyridine displays sizable contributions from an partial wave, which can be identified as being weakly trapped from the very low amplitude of its quasi-bound part, the contributions of are much weaker in pyrimidine and barely visible in -triazine as a consequence of symmetry.
-The three-fold (D) symmetry of the -triazine scaffold is indeed the reason why and partial waves are dominant and other partial waves are suppressed in that case.
-The suppression of these lower-, more weakly trapped partial waves, to which the partial waves can couple to release the photoelectron more efficiently, explains the increasing lifetime of the shape resonance in the series pyridine, pyrimidine, -triazine.
+In addition to the changing probability density ratio between the wavefunctions in the quasi-bound and tunneling regions, we also observe the manifestation of molecular symmetry.
+The reason that the partial wave prevails in the resonant wavefunctions can be qualitatively understood from the inset panel in Fig.
+5H.
+The resonant wavefunction displays five nodal surfaces, as qualitatively illustrated in the inset of Fig.
+5H, corresponding to a dominant partial-wave character.
+However, the wavefunction also contains contributions from the other partial waves.
+As the symmetry of the molecular scaffold of the neutral molecule decreases from -triazine (D3h) to pyrimidine and pyridine (C2v), the relative amplitudes of to partial waves increases as a consequence of lower symmetry.
+The increasing relative amplitudes of the untrapped partial waves cause a faster decay of the resonant state.
 To summarize, these two mechanisms explain the trend in the resonant photoelectron trapping times in the azabenzene series.
@@
 5D,F,H, which show that the photoionization delays in the vicinity of the discussed shape resonance indeed increase in the sequence pyridine, pyrimidine, -triazine.
+V Conclusion
 In conclusion, we have used the novel attosecond capabilities of LCLS to measure core-level photoionization delays of aromatic molecules.
 These measurements were made possible by the combination of attosecond X-ray pulses with the angular streaking technique.
-Absolute N-1s photoionization delays were accessed by using the much faster C-1s photoelectrons as a time reference.
+Absolute N-1s photoionization delays were accessed by using the much faster and therefore prompt C-1s photoelectrons as a time reference.
 These self-referenced measurements were realized on a fine energy grid and revealed the presence of local maxima at 3-15 eV above the N-1s threshold.
@@
 The two dominant mechanisms governing the photoelectron trapping were identified as (i) changes in the barrier height of the molecular potential that lead to an increase of the electron trapping with increasing electronegativity of the functional groups and (ii) the increasing symmetry of the molecular scaffold reduces the coupling between the well-trapped high- and the less-trapped lower- states.
-Both effects individually and collectively increase the trapping time in the sequence pyridine, pyrimidine, -triazine.
+Both effects individually and collectively increase the trapping time in -triazine compared to pyridine.
 These results demonstrate the promise of performing attosecond core-level photoemission experiments at free-electron-laser facilities for the following reasons.
-The studied molecules are indeed key building blocks of biologically active molecules, as well as molecules employed in optoelectronics, which opens the perspective of studying attosecond electron dynamics in these systems.
-The key advantage of core-level photoionization delays over valence-shell photoionization delays, is that the photoelectron spectra always remain simple, featuring only one dominant peak per atom of a given type.
-This will enable the application of the methods described in this work to complex molecules, molecular assemblies and polymers, such as organic semiconductors.
-Specifically, the effects governing the attosecond electron trapping in aromatic systems unveiled in the present work may become guiding principles for the design and interpretation of experiments studying charge migration and charge transfer in functional materials.
+The studied molecules are indeed key building blocks of biologically active molecules, as well as molecules employed in optoelectronics, which opens the perspective of extending attosecond chronoscopy to these systems.
+The key advantage of core-level photoionization delays over valence-shell photoionization delays, is that the photoelectron spectra always remain comparably simple.
+This will enable the application of the methods described in this work to complex molecules, molecular assemblies and polymers, such as organic semiconductors, which offers the perspective of studying attosecond electron scattering and trapping in functional materials.
+VI Methods
+Experimental methodology
+The attosecond soft X-ray pulses were produced using the photocathode shaping method [40].
+The mean pulse energy is J and the mean FWHM bandwidth is eV.
+From our previous measurements, we can estimate a FWHM temporal pulse duration of as for the X-ray pulses.
+The attosecond soft X-ray pulses described above were focused to a spot-size of m in diameter using a pair of Kirkpatrick-Baez focusing mirrors [41].
+The -nm streaking pulse with a duration of 40 fs was produced by a commercial optical parametric amplifier (TOPAS-HE, Light Conversion) which is pumped by a fs, mJ, -nm laser source.
+The signal beam was sent through a superachromatic quarter-wave plate (Thorlabs SAQWP05M-1700) to produce a circularly polarized laser field and was focused with a mm CaF2 lens before being reflected from a silver-coated mirror with a -mm-diameter hole to co-propagate the laser and X-ray fields to the interaction point in the c-VMI spectrometer chamber.
+High pressure (6.4 bar) helium gas was bubbled through a vessel heated to 80-140∘C to deliver the samples into the c-VMI spectrometer chamber, and the photoelectrons ionized from the samples were collected using an electrostatic lens.
+A variable-line spaced (VLS) grating X-ray photon spectrometer [42, 43] was employed downstream of the interaction point to obtain the spectra for individual X-ray shots.
+Data analysis
+The relative photoionization delay between the carbon and nitrogen K-shell photoemission features, , is related to the difference in streaking angle, , between the momentum shift of the nitrogen and carbon photoemission features, , where fs is the period of the IR streaking laser.
+We use two methods to extract the differential streaking angle from the dataset.
+The first method uses a partial covariance approach to isolate the correlated changes of streaking-induced variations in the photoelectron momentum distribution, while the second procedure uses a sorting method which tags the arrival time of the attosecond X-ray pulse relative to the IR-field.
+The two approaches are further illustrated and compared in the supplementary information.
+Theoretical methodology
+The N photoionization of the three azabenzenes considered here were studied using two methods, the basis-set complex Kohn method [44] and ePolyScat [45, 46].
+The one-electron basis set used to obtain the occupied molecular orbitals was a correlation consistent valence triple zeta [47] and the geometry used minimized the Hartree-Fock energy.
+The ePolyScat calculations used a static-exchange potential with equivalent-core method.
+An additional CP potential was introduced to approximate the interaction between the emitted electron and the residual ion of the molecule [48].
+The calculated values with and without the CP potential are compared in the SM.
+The multichannel calculation from the coupled N holes using the complex-Kohn method [49] and the one-channel calculation using ePolyScat from the localized N hole show consistency, as demonstrated in the SM.
+Data availability
+All data will be made available on the public ETH data repository upon acceptance of the manuscript.
 Acknowledgments
-Funding: Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S.
+Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S.
 Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No.
 DE-AC02-76SF00515.
-The work of J.-B.J., X.G., D.H., M.H., S.H.
+The work of J.-B.J., X.G., D.H., M.H.
+(ETH Zürich), S.H.
 and H.J.W.
@@
 The work of M.H.
-was additionally supported by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 801459 - FP-RESOMUS.
-The effort from T.D.D., J.T.O.
-A.L.W., J.W., T.J.A.W and J.P.C.
+(ETH Zürich) was additionally supported by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 801459 - FP-RESOMUS.
+The effort from T.D., J.T.O.
+A.L.W., J.W., T.J.A.W.
+and J.P.C.
 was supported by DOE, BES, Chemical Sciences, Geosciences, and Biosciences Division (CSGB).
-Z.G, D.C., J.D., P.L.F., R.R.
+Z.G., D.C., J.D., P.L.F., R.R.R.
 and A.M.
@@
 Z.G., P.L.F.
-and R.R.
+and R.R.R.
 also acknowledge support from Robert Siemann Fellowship of Stanford University.
@@
 was supported in part by the Berkeley Lab Undergraduate Faculty Fellowship (BLUFF) Program, managed by Workforce Development & Education at Berkeley Lab.
-Author Contributions: J.-B.J., T.D., K.U., A.M., J.P.C.
+Author contributions
+J.-B.J., T.D., K.U., A.M., J.P.C.
 and H.J.W.
@@
 wrote the initial draft of the manuscript, which was reviewed and edited by all authors.
-Competing Interests: The authors declare no competing interests.
-Data availability statement: All data will be made available on the public ETH data repository upon acceptance of the manuscript.
+Competing interests
+The authors declare no competing interests.